OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the given curves about the specified line. Y=x^2, x=y^2; about y =1

OpenStudy (dumbcow):

\[V =\pi \int\limits_{0}^{1}(1-x^{2})^{2}-(1-\sqrt{x})^{2} dx\]